OpenStudy (anonymous):

if dy/dx=4y and y=4 when x=0, then y= a)4e^4x b)e^4x c)3+e^4x d)4+e^4x e)2x^2+4

OpenStudy (anonymous):

I feel like something is missing in this question.

OpenStudy (anonymous):

Never mind.

jimthompson5910 (jim_thompson5910):

dy/dx=4y dy=4y*dx dy/y = 4dx int(dy/y) = int(4dx) ln|y| = 4x+C y = +-e^(4x+C) y = +-e^(4x)*e^C y = +-e^C*e^(4x) y = A*e^4x From here, plug in x = 0 and y = 4, then solve for A